# Ev2R: Evaluating Evidence Retrieval in Automated Fact-Checking

### v1

@@ 5 Results and Discussion
The results indicate that the choice of backbone LLMs does impact the performance of prompt Ev2R scorers to some extent, highlighting the importance of selecting appropriate models for optimal scoring performance. However, no single backbone model consistently outperforms its counterparts across all dimensions. For example, while the Gemini-based scorers outperforms in the coverage dimension, the GPT-$4o$-based scorer excels in the category of verdict agreement.
Scorer Coverage Coherence Repetition Consistency Relevance Verdict Agree Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360 .240 .243 .179 .205
Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313 .278 .256 .198 .197
Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269 .314 .284 .196 .195
Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221 .523 .483 .207 .223
Prompt Scorer Gemini-Pro
Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263 .262 .273 .177 .200
Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264 .237 .177 .183 .157
Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248 .285 .278 .207 .203
Prompt Scorer Gemini-Flash
Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333 .230 .247 .204 .217
Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368 -.290 -.205 .041 .020
Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531 -.190 -.039 .054 .090
Prompt Scorer Llama 3.1
Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250 .215 .208 .150 .197
Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172 .216 .210 .138 .148
Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187 .128 .121 .144 .149
Trained Scorer
Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110 .151 .098 .055 .063
Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374 .487 .459 .265 .275
Baselines
RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099 .070 .097 .103 .119
BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079 .166 .137 .094 .072
Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076 .152 .159 .085 .100
H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003 -.029 -.019 .034 .003
Table 2: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers. We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients.
Comparing the scores obtained from the AVeriTeC test set evaluation (Table 1) with those from the shared task submissions (Table 2) reveals that some scorers, such as the reference-based (recall) ones, maintain consistent performance across datasets, while others do not. This difference in performance can be partially attributed to the distinct specifications of the two datasets. While the AVeriTeC subset (Table 1) was scored by two or three human evaluators (in cases of disagreement), it includes only evidence retrieved by a single system - the AVeriTeC baseline system. In contrast, the shared task predictions were extracted from all systems submitted to the AVeriTeC shared task.
@@
In Table 3, we assess the performance of Ev2R scorers across various adversarial tests. In most test categories, both the prompt-based reference scorers and the trained proxy-reference scorers demonstrate strong robustness. While test categories that introduce limited noise into the evidence data (e.g., typos or redundant words) lead to significant performance drops in baseline scorers and the trained reference-based scorer, they have little to no impact on the prompt-based and trained proxy-reference scorers. Results indicate that most of the proposed Ev2R scorers can handle linguistic variations and redundancy in evidence better than the baselines.
5.1 Qualitative analysis of scorers
Through manual assessment of evidence evaluation by different scorers, we observe various strengths and limitations across reference-less, reference-based, and proxy-reference approaches.
Reference-less
The reference-less Ev2R scorers evaluate supporting or refuting evidence accurately, but struggle with cases where there is not enough evidence as they lack any reference (i.e., evidence or verdict label) to compare the predicted evidence against. For example, for the claim “France24 encouraging Niger Delta militants to fight for their right to equity in resource sharing.” (id 20) in Table 8 in the appendix, the predicted evidence does not provide any supporting or refuting information. While it discusses groups in the Niger Delta, no information is given on whether France24 encourages them or not. Moreover, for some claims, unclear formulations pose challenges for reference-less scorers when extracting facts from the claims to evaluate against the predicted evidence. For instance, in claim “Antifa vandalized the home of Missouri senator Josh Hawley in 4th of January 2021” (593), it is unclear if ‘Antifa’ refers to a specific organisation, a left-wing coalition of organisations, or a specific event such as a demonstration. Finally, reference-less scoring struggles with extensive evidence that includes both supporting and refuting information. On the other hand, breaking down the claim into multiple facts enhances understanding of more complex claims and their evaluation process. An example is claim “Biden won fewer counties than Trump and Obama but more votes, suggesting election fraud in 2020.” (648), which is decomposed into five facts (see Table 8 in the appendix).
Dataset METEOR ROUGE BLEU BLEURT Ref-less Prompt ref-based Trained ref-based Proxy-ref
Prec Rec
Completeness -30.56 -53.19 -77.87 -143.70 -91.76 -21.2 -59.5 -82.68 -39.35
Random shuffle -32.32 -89.94 -94.20 -1.02 0.55 -6.8 -6.0 -92.32 -32.22
Average -31.44 -71.57 -86.04 -72.36 -45.61 -14.0 -32.75 -87.50 -35.79
Invariance contraction -0.53 -0.97 -1.35 -1.65 0 -0.8 -0.7 0.02 -0.85
Invariance num2text -8.34 -10.24 -14.10 -22.13 -3.85 0.1 0 -35.64 -4.02
Invariance text2num -0.04 -0.90 -1.15 -0.81 -4.40 0 -0.3 0.01 -1.59
Invariance synonyms -8.22 -36.64 -51.83 -51.00 -2.75 -11.6 -11.3 -85.71 -15.35
Redundancy sent -26.16 -46.52 -56.15 -80.45 3.30 0 -0.1 -67.95 -5.04
Redundancy words -5.73 -10.68 -32.69 -31.90 -2.75 -1.0 0.2 -52.55 -2.89
Fluency -11.52 -23.33 -29.26 -41.69 -2.20 -4.3 -3.9 -57.12 -2.6
Noise -31.82 -31.81 -31.81 -21.20 -1.65 -28.7 -2.2 -62.42 2.89
Argument structure -0.01 0 0 0 1.65 0 0 -28.93 -0.74
Average -10.33 -18.43 -23.94 -27.87 -15.04 -5.16 -2.9 -39.79 -3.78
Table 3: Results obtained for adversarial checklist tests as score difference (in %) between initial evidence and manipulated evidence. The evaluated prompt and trained Ev2R scorers use $GPT4o$ and DeBERTa as base models, respectively. Overall, the prompt reference-based scorers and the proxy-reference scorers demonstrate the most robust performance on adversarial tests, as evidenced by the small performance drops.
@@ 5.1 Qualitative analysis of scorers
The reference-based approach allows for a more detailed evaluation by breaking down both reference and predicted evidence into atomic facts (see Table 5 in the appendix). Additionally, reference evidence provides extra context to the Ev2R scorer, which can be leveraged during the evaluation of retrieved evidence. Moreover, communicating precision and recall through distinct scores also enables a more detailed understanding of the predicted evidence. For instance, retrieved evidence may sometimes include excessive details compared to the reference, affecting precision but still yielding high recall scores While this approach enhances the evaluation process, we observed some limitations during qualitative analysis. One of the main challenges with reference-based scoring is that the predicted evidence often uses different information or reasoning chains than the reference evidence. Even when both reference and predicted evidence lead to the same conclusion (e.g., the claim is refuted as in Table 5 in the appendix), the score is generally lower compared to reference-less and proxy-reference scorers. For example, in Table 5 both pieces of evidence contradict the claim that Nigeria is the fifth largest recipient of diaspora remittances. However, the predicted evidence states that Nigeria is the sixth largest recipient, while the reference states it is the seventh largest. As a result, the scorer finds that the reference and predicted evidence do not fully align. Furthermore, retrieved evidence can sometimes include more details than the reference, which impacts precision but leads to a high recall.
@@
Our analysis further shows that the proxy-reference Ev2R scorer often yields lower scores due to differences between predicted and reference evidence, particularly when one supports and the other refutes the given claim. For example, in Table 7 in the appendix, the predicted evidence supports the claim, while reference evidence does not. On the other hand, when the predicted evidence uses different information or reasoning approaches than the reference evidence but still reaches the same verdict (i.e., the proxy reference), the score tends to be higher compared to scores obtained from the reference-based scorers. However, minor numerical differences, such as in ID 56, can still result in diverging verdicts despite overall consistency between the reference and predicted evidence.
Overall, across all scores, when comparing the Ev2R scorers to the evaluated baselines, we find that repeated information in the predicted evidence and linguistic variations do not impact the scores, as the evaluation remains robust against these aspects. For example, in Table 6 in the appendix, the predicted evidence contains multiple repetitions and is significantly longer than the reference evidence. This results in relatively low METEOR and ROUGE scores of $0.26$ and $0.11$, while the reference-less, reference-based, and proxy reference scorer return the highest possible score of $1$.
@@ 6 Conclusion
This paper introduced an evidence evaluation framework for AFC, proposing three types of evidence scorers: reference-less, reference-based, and proxy-reference evaluation. We evaluate the framework using the AVeriTeC dataset, including evidence retrieved by the AVeriTeC baseline system as well as submitted systems from the AVeriTeC shared task. We assess the proposed Ev2R scorers based on human ratings across multiple dimensions (e.g., coverage and relevance) and by testing their sensitivity to automated perturbations in the evidence data. Our results show that the reference-based atomic Ev2R scorers correlate best with human ratings and are robust to automated perturbations. However, they struggle with evidence that uses different information or reasoning chains than the reference evidence—a scenario where LLM-based reference-less and proxy-reference scorers perform better. Overall, our proposed Ev2R scorers excel compared to traditional evaluation metrics like METEOR and ROUGE.
@@ Acknowledgements
We thank Nedjma Ousidhoum, Julian Eisenschlos, Oana Cocarascu, and Elena Simperl for their valuable feedback and suggestions about this work. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958). Michael Schlichtkrull is supported by the EPSRC grant AdSolve (EP/Y009800/1), a RAI UK Keystone project.
Appendix A Adversarial checklist tests
@@
Appendix B Prompt templates

### v2

@@ 5 Results and Discussion
Assessing the impact of different backbone LLMs for the reference-based component of Ev2R, we find that the model selection has only a limited effect on the overall performance of the Ev2R scorer across datasets (Table 2) and evaluation dimensions (Table 2 and 4). No single model consistently outperforms others across all dimensions and datasets. While scorer variants using Gemini models correlate higher with human assessments in the dimensions coverage and relevance, the GPT-$4o$-based variant excels in the category of verdict agreement (Table 2). Moreover, the combination of GPT-$4o$ and proxy-reference shows strong correlations on AVeriTeC and is competitive overall.
@@
In Table 5, we assess the performance of Ev2R components (reference-based and proxy-reference) across various adversarial tests against baselines, considering both semantics-preserving and semantics-altering tests. Our insights highlight how Ev2R components (i.e., reference-based and proxy-reference) versus baseline scorers respond to adversarial manipulations in evidence texts. Traditional metrics such as METEOR, ROUGE, BLEU, and especially BLEURT show high sensitivity to semantics-altering tests, demonstrated through substantial performance drops. On the other hand, the Ev2R components, specifically proxy-reference scoring, exhibit stronger robustness to both semantics-altering and semantics-preserving evaluation. These findings highlight their suitability and robustness for evidence evaluation. Overall, Ev2R demonstrates greater robustness compared to baselines in adversarial settings.
5.1 Qualitative analysis of evidence evaluation
@@
Reference evidence provides additional context to the Ev2R scorer, which can be leveraged during the evaluation of retrieved evidence. The reference-based scoring component of Ev2R allows a more detailed and interpretable evaluation by breaking down both reference and predicted evidence into atomic facts (see Table 9 in the appendix). Reporting separate precision and recall scores facilitates a more detailed understanding of the retrieved evidence. For instance, retrieved evidence sometimes contains excessive details compared to the reference, negatively affecting precision but still yielding high recall.
While this approach enhances the evaluation process, our qualitative analysis identified challenges that are addressed by stabilizing the Ev2R scorer through integration of a proxy-reference component. One specific challenge for reference-based scoring are cases where predicted evidence uses different information or follows a different reasoning chain than the reference evidence, but results in the same verdict (e.g., refuting a claim, as illustrated in Table 9 in the appendix). In such cases, the reference-based scoring tends to yield lower scores compared to reference-less and proxy-reference scorers. For example, in Table 9 both reference and predicted evidence contradict the claim that Nigeria is the fifth largest recipient of diaspora remittances. However, the predicted evidence states that Nigeria is the sixth largest recipient, while the reference states it is the seventh largest. Hence, the scorer identifies a mismatch between the reference and predicted evidence. Furthermore, retrieved evidence can sometimes include more details than the reference, which impacts precision but leads to a high recall. To address these limitations and stabilise the Ev2R scorer, we incorporate proxy-reference scoring.
@@
We also observe that a key challenge for proxy-reference scoring are evidence pairs where both predicted and reference evidence are related but result in diverging verdict labels (see Table 11 in the appendix). For example, considering the claim “Scotland is spending more on health per head than the rest of the UK.” in Table 11; while the predicted evidence closely resembles the reference in surface-level characteristics, the predicted evidence supports the claim, whereas the reference evidence does not, resulting in a low proxy-reference score. Similarly, minor numerical differences, such as those observed in evidence in Table 11, can also lead to diverging verdicts despite overall consistency between the reference and predicted evidence. In these cases, a more interpretable assessment based on precision and recall scores from reference-based scoring can yield further insights and enhance readers’ understanding of the presented evidence.
5.2 How Ev2R addresses key challenges of evidence evaluation for AFC
This section reiterates key challenges of evidence evaluation (see Sections 1 and 2) and Ev2R.
Implicit evidence assessment
Verdict-based evaluation approaches can inaccurately rate evidence high by predicting correct verdicts without genuinely relying on the retrieved evidence. To overcome this over-reliance on verdict alone, Ev2R explicitly evaluates both alignment with reference evidence and verdict agreement. Evaluation results from AVeriTeC, FEVER, and VitaminC show high correlation with human judgments on criteria such as verdict agreement, evidence relevance, and coverage of reference data, indicating Ev2R assesses retrieved evidence rather than verdict correctness alone.
Assessment relying on closed knowledge sources & surface-level matching of evidence
Previous widely-used approaches for evidence evaluation (e.g., the FEVER score (Thorne et al., 2018a)) relied on predefined sources (e.g., Wikipedia) and exact matches between predicted and reference evidence. Such methods are not extendable to realistic AFC settings, where predicted evidence may not be part of the annotated corpus. Additionally, token-matching metrics (e.g., h-METEOR used in (Schlichtkrull et al., 2024b)) are overly sensitive to surface-level differences and fail to capture alternative valid evidence or reasoning paths. Ev2R’s reference-based component decomposes retrieved evidence into atomic facts and assesses their alignment with reference facts beyond strict matching constraints. Evaluations of evidence predicted with shared task systems (see Table 4), which include unannotated web evidence, confirm the effectiveness of this approach. Furthermore, adversarial test results (see Table 5) demonstrate Ev2R’s superior robustness against surface-level perturbations compared to baseline scorers.
Multiple valid evidence sets and reasoning paths & temporal constraints with evidence annotation
For a given claim, multiple correct sets of evidence or alternative reasoning paths can exist. Previous evaluation methods might penalize valid alternative evidence. Ev2R with its both scoring compnents considers scenarios with multiple valid reasoning paths and evidence absent in the annotated references. Evaluations using the datasets FEVER, VitaminC, and AVeriTeC, where exact matching between reference and predicted evidence is not expected, show Ev2R achieves higher correlations with human judgments and stable performance under adversarial testing.
@@ 6 Conclusion
This paper introduces Ev2R, a weighted evidence evaluation scorer for AFC. Ev2R consists of a reference-based and a verdict-level proxy evaluation component and addresses limitations in previous evaluation approaches by jointly assessing how well evidence aligns with gold references and how reliably the evidence supports the verdict. We assessed Ev2R against three baseline approaches for evidence evaluation approaches and find that Ev2R correlates strongly with human judgments on relevant evaluation criteria and is a robust scorer for evidence evaluation in AFC.
7 Limitations & Ethics statement
This work has several limitations that are important to acknowledge. First, the datasets used are restricted to English, limiting the generalisability of findings to global fact-checking contexts where multilingual capabilities are essential. This English-only focus reflects a broader issue in the field, where most benchmarks are available solely in English, potentially distorting perceptions of progress in automated fact-checking. Additionally, all labels, e.g., supports and refutes used in this work to classify claims, capture the relationship between claims and evidence, they do not imply any real-world correctness of the claims. Finally, We informed the participants about the data being collected and its purpose. Participants had the opportunity to withdraw at any time and to provide feedback.
@@ Acknowledgements
We thank Nedjma Ousidhoum, Julian Eisenschlos, Oana Cocarascu, and Elena Simperl for their valuable feedback and suggestions about this work. Mubashara Akhtar acknowledges funding from the ETH AI Center Postdoctoral Fellowship. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958). Michael Schlichtkrull is supported by the EPSRC grant AdSolve (EP/Y009800/1), a RAI UK Keystone project.
Appendix A Evaluation with adversarial perturbations
@@
Appendix B Additional Evaluation Dimensions
As an additional data source for the evaluation of scorers, we collected ratings for 100 randomly selected test data instances from the AVeriTeC test set (Schlichtkrull et al., 2023), together with the evidence retrieved through the AVeriTeC baseline system as described in Schlichtkrull et al. (2023). Each test sample was evaluated by two annotators, who were either computer science graduate students or postdoctoral researchers familiar with AFC research. This resulted in 200 annotations while the overall rating for each annotated sample and dimension was calculated as the average of both annotations. If the annotators disagreed on the verdict assigned to the retrieved evidence, hence interpreted the relationship between retrieved evidence and claims differently, we assigned the data instance to a third annotator and used majority voting to determine the final verdict.
Scorer Coverage Coherence Repetition Consistency Relevance
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360
Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313
Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269
Ref-based ($F_{1}$) .306 .298 .192 .194 -.109 -.097 .201 .213 .295 .289
Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221
Prompt Scorer Gemini-Pro
Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263
Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264
Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248
Ref-based (F1) .345 .304 .224 .203 -.120 -.135 .196 .201 .259 .256
Prompt Scorer Gemini-Flash
Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333
Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368
Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531
Ref-based (F1) .104 .249 .104 -.008 -.524 -.521 .333 .223 .508 .435
Prompt Scorer Llama 3.1
Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250
Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172
Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187
Ref-based (F1) .318 .324 .176 .180 -.080 -.069 .104 .125 .150 .179
Trained Scorer
Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110
Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374
Baselines
RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099
BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079
Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076
H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003
Weighted Scorer ($\alpha=0.5$)
GPT-4o / Proxy-ref .321 .323 .211 .240 -.060 -.077 .247 .271 .297 .332
Gemini-Pro / Proxy-ref .341 .326 .227 .244 -.066 -.096 .245 .265 .278 .315
Gemini-Flash / Proxy-ref .221 .299 .167 .147 -.267 -.289 .313 .276 .403 .404
Llama 3.1 / Proxy-ref .328 .336 .203 .233 -.046 -.063 .199 .227 .224 .277
Table 7: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers, across five quality dimensions. Highest scores per column are colored blue and second-highest values highlighted with brown color. Reference-based precision and recall scores which are calculated to compute the final reference-based $F_{1}$ score, are colored in gray.
Table Table 8 given an overview over correlation between scorers and human ratings on three additional dimensions, shortly outlined below.
$(1)$ Coherence captures whether the retrieved evidence is coherent, i.e., all sentences are connected sensibly and the evidence makes sense as a whole. An important aspect of fact-checking is presenting the verification process to the reader in a way that is comprehensible. Hence, well-structured and coherent evidence supports understanding of the verification process and the rationale behind the resulting verdict.
$(2)$ Repetition evaluates whether the retrieved evidence exhibits repetition of its content. As mentioned earlier, supporting readers’ understanding of the fact-checking process can be enhanced through avoiding repetitive content in the evidence. This can further maintain the reader’s focus on unique information in the evidence text, leading to better understanding of the resulting conclusion, i.e., the verdict.
$(3)$ Consistency assesses whether the retrieved evidence is logically consistent in itself, since consistency in evidence improves the reliability of the fact-checking process and trust among readers.
Scorer Coverage Coherence Repetition Consistency Relevance
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360
Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313
Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269
Ref-based ($F_{1}$) .306 .298 .192 .194 -.109 -.097 .201 .213 .295 .289
Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221
Prompt Scorer Gemini-Pro
Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263
Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264
Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248
Ref-based (F1) .345 .304 .224 .203 -.120 -.135 .196 .201 .259 .256
Prompt Scorer Gemini-Flash
Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333
Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368
Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531
Ref-based (F1) .104 .249 .104 -.008 -.524 -.521 .333 .223 .508 .435
Prompt Scorer Llama 3.1
Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250
Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172
Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187
Ref-based (F1) .318 .324 .176 .180 -.080 -.069 .104 .125 .150 .179
Trained Scorer
Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110
Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374
Baselines
RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099
BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079
Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076
H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003
Weighted Scorer ($\alpha=0.5$)
GPT-4o / Proxy-ref .321 .323 .211 .240 -.060 -.077 .247 .271 .297 .332
Gemini-Pro / Proxy-ref .341 .326 .227 .244 -.066 -.096 .245 .265 .278 .315
Gemini-Flash / Proxy-ref .221 .299 .167 .147 -.267 -.289 .313 .276 .403 .404
Llama 3.1 / Proxy-ref .328 .336 .203 .233 -.046 -.063 .199 .227 .224 .277
Table 8: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers, across five quality dimensions. Highest scores per column are colored blue and second-highest values highlighted with brown color. Reference-based precision and recall scores which are calculated to compute the final reference-based $F_{1}$ score, are colored in gray.